# On the Brun–Titchmarsh theorem

### v1

@@ 5.2. Approaching by multiplicative characters
In this subsection, we assume that $q$ is squarefree and $\eta$-smooth, as required in Theorem 1.5. The application of Lemma 3.2 will depend on the sizes of corresponding Dirichlet polynomials, we divide the non-trivial characters $\,(\operatorname{mod}{q})$ into eight sets. Precisely, we define
$\Gamma_{1}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{1}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\}.$
@@
$S_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)L(s,\chi)|,$
$T_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{W}(s,\chi)\mathcal{L}(s,\chi)|.$
@@
For $\Gamma\in\{\Gamma_{2}^{+},\Gamma_{4}^{-},\Gamma_{4}^{+},\Gamma_{4}^{-}\},$ we have
@@
(5.18) $\sum_{\chi\in\Gamma_{k}^{+}}|\mathcal{W}\mathcal{L}|^{4}\ll(WL)^{2}q^{\varepsilon}$
(5.19) $\sum_{\chi\in\Gamma_{k}^{-}}|\mathcal{W}\mathcal{L}|^{12}\ll(WL)^{2}q^{1+\varepsilon}.$
@@
We also assume that
(5.22) $W\leqslant q^{\frac{\delta}{2}},$
so that Conjecture 1.1 yields
(5.23) $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|\mathcal{W}(s,\chi)L(s,\chi)|^{4}\ll q^{1+\varepsilon}.$
(I) Treatment to $T_{1}^{\pm}$: From (5.21) and the trivial bound $\mathcal{W}\ll W^{\frac{1}{2}}$, we obtain
$T_{1}^{\pm}\ll|s|(WL)^{\frac{1}{2}-}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}||\mathcal{N}|.$
By Cauchy–Schwarz inequality and Lemma 5.1, we have
$T_{1}^{\pm}\ll|s|(WL)^{\frac{1}{2}-}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{2}}\ll|s|x^{\frac{1}{2}-}$
@@
$S_{2}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{\pm}}|\mathcal{W}L|^{4}\Big{)}^{\frac{1}{4}}.$
Hence Lemma 5.1, (5.17) and (5.23) yield
@@
Using (5.18) or (5.19) in place of (5.23), we obtain
$T_{2}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{+}}|\mathcal{W}\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{1}{8}}(Nq)^{\frac{1}{8}}(W^{2}L^{2})^{\frac{1}{4}}$
$\ll q^{\frac{1}{4}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}(WL)^{\frac{1}{2}}$
@@
$T_{2}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{3}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{24}}\Big{(}\sum_{\chi\in\Gamma_{2}^{-}}|\mathcal{W}\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{3}{8}}(Nq)^{\frac{1}{24}}(qW^{2}L^{2})^{\frac{1}{12}}$
$\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{24}}(WL)^{\frac{1}{6}}.$
@@
$\min\{S_{2}^{+},~{}T_{2}^{+}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,q^{-\frac{1}{4}}(WL)^{\frac{1}{2}}\}\ll x^{\frac{1}{8}}q^{\frac{7}{16}}M^{\frac{3}{8}}q^{\varepsilon}$
@@
$\min\{S_{2}^{-},~{}T_{2}^{-}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,N^{-\frac{1}{12}}(WL)^{\frac{1}{6}}\}\ll x^{\frac{1}{12}}q^{\frac{1}{2}}M^{\frac{5}{12}}q^{\varepsilon}.$
@@
$S_{4}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\chi\in\Gamma_{4}^{+}}|\mathcal{W}L|^{4}\Big{)}^{\frac{1}{4}}.$
From Lemma 5.1, (5.16), (5.17) and (5.23), it follows that
@@
Using (5.18) or (5.19) in place of (5.23), we obtain
$T_{4}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{W}\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}q^{\frac{5}{16}}q^{\frac{5}{16}}(Mq)^{\frac{1}{6}}(Nq)^{\frac{1}{6}}(W^{2}L^{2})^{\frac{1}{4}}$
$\ll q^{\frac{3}{4}+\varepsilon}(MN)^{\frac{1}{16}}(WL)^{\frac{1}{2}}$
@@
$T_{4}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{W}\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}q^{\frac{7}{16}}q^{\frac{7}{16}}(Mq)^{\frac{1}{48}}(Nq)^{\frac{1}{48}}(qW^{2}L^{2})^{\frac{1}{12}}$
$\ll q^{1+\varepsilon}(MN)^{\frac{1}{48}}(WL)^{\frac{1}{6}}.$
@@
$\min\{S_{4}^{+},~{}T_{4}^{+}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,q^{-\frac{1}{4}}(WL)^{\frac{1}{2}}\}\ll x^{\frac{1}{16}}q^{\frac{31}{32}+\varepsilon}$
@@
$\min\{S_{4}^{-},~{}T_{4}^{-}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,(WL)^{\frac{1}{6}}(MN)^{-\frac{1}{24}}\}\ll x^{\frac{1}{20}}q^{1+\varepsilon}.$
@@
Combining the above treatments in four parts, as well as the restrictions (5.20), (5.24), (5.25) and (5.22), we choose
$M=N=x^{1-\varepsilon}q^{-\frac{6}{5}},\ \ W=\min\{x^{1-\varepsilon}(MN)^{-1},~{}q^{\frac{\delta}{2}}\}$
@@
$D=\min\{x^{1-\varepsilon},~{}x^{2-\varepsilon}q^{\frac{\delta}{2}-\frac{12}{5}}\}$
@@
Keep in mind that $q\geqslant x^{\frac{1}{2}}$ throughout this section.

### v3

@@ 5.1. Approaching by multiplicative characters
In this subsection, we assume $q$ is squarefree and $\eta$-smooth, as required in Theorem 1.5. The application of Lemma 2.2 will depend on the sizes of corresponding Dirichlet polynomials, so that we divide all non-trivial characters $\,(\operatorname{mod}{q})$ into eight sets. Precisely, we define
$\Gamma_{1}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{1}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{2}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{3}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\geqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{+}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\geqslant q^{\frac{1}{8}}\},$
$\Gamma_{4}^{-}=\{\chi\neq\chi_{0}\,(\operatorname{mod}{q}):|\mathcal{M}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{N}(s,\chi)|\leqslant q^{\frac{1}{4}},|\mathcal{L}(s,\chi)|\leqslant q^{\frac{1}{8}}\}.$
@@
$S_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)L(s,\chi)|,$
$T_{k}^{\pm}$ $=\sum_{\begin{subarray}{c}\chi\in\Gamma_{k}^{\pm}\end{subarray}}|\mathcal{M}(s,\chi)\mathcal{N}(s,\chi)\mathcal{L}(s,\chi)|.$
@@
For $\Gamma\in\{\Gamma_{2}^{+},\Gamma_{2}^{-},\Gamma_{4}^{+},\Gamma_{4}^{-}\},$ we have
@@
(5.16) $\sum_{\chi\in\Gamma_{k}^{+}}|\mathcal{L}|^{4}\ll L^{2}q^{\varepsilon}$
(5.17) $\sum_{\chi\in\Gamma_{k}^{-}}|\mathcal{L}|^{12}\ll L^{2}q^{1+\varepsilon}.$
@@
We also note that
$\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|L(s,\chi)|^{4}\ll q^{1+\varepsilon}.$
(I) Treatment to $T_{1}^{\pm}$: From (5.19) it follows that
$T_{1}^{\pm}\ll|s|L^{\frac{1}{2}-}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}||\mathcal{N}|.$
By Cauchy–Schwarz inequality and (5.12), (5.13), we have
$T_{1}^{\pm}\ll|s|L^{\frac{1}{2}-}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{1}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{2}}\ll|s|x^{\frac{1}{2}-}$
@@
$S_{2}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{\pm}}|L|^{4}\Big{)}^{\frac{1}{4}}.$
Hence Lemma 5.1 and (5.15) yield
@@
Using (5.16) and (5.17), we obtain
$T_{2}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{8}}\Big{(}\sum_{\chi\in\Gamma_{2}^{+}}|\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{1}{8}}(Nq)^{\frac{1}{8}}L^{\frac{1}{2}}$
$\ll q^{\frac{1}{4}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}L^{\frac{1}{2}}$
@@
$T_{2}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{3}{8}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{2}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{24}}\Big{(}\sum_{\chi\in\Gamma_{2}^{-}}|\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}M^{\frac{1}{2}}q^{\frac{3}{8}}(Nq)^{\frac{1}{24}}(qL^{2})^{\frac{1}{12}}$
$\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{24}}L^{\frac{1}{6}}.$
@@
$\min\{S_{2}^{+},~{}T_{2}^{+}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,q^{-\frac{1}{4}}L^{\frac{1}{2}}\}\ll x^{\frac{1}{8}}q^{\frac{7}{16}}M^{\frac{3}{8}}q^{\varepsilon}$
@@
$\min\{S_{2}^{-},~{}T_{2}^{-}\}\ll q^{\frac{1}{2}+\varepsilon}M^{\frac{1}{2}}N^{\frac{1}{8}}\min\{1,N^{-\frac{1}{12}}L^{\frac{1}{6}}\}\ll x^{\frac{1}{12}}q^{\frac{1}{2}}M^{\frac{5}{12}}q^{\varepsilon}.$
@@
In fact, the restriction $q\leqslant x^{\frac{5}{11}-\varepsilon}$ is additionally imposed to guarantee that $M$ is allowed to be bigger than $q$.
@@
$S_{4}^{\pm}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{\pm}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\chi\in\Gamma_{4}^{+}}|L|^{4}\Big{)}^{\frac{1}{4}}.$
From Lemma 5.1, (5.14) and (5.15), it follows that
@@
Using (5.16) and (5.17), we obtain
$T_{4}^{+}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{5}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{+}\end{subarray}}|\mathcal{L}|^{4}\Big{)}^{\frac{1}{4}}$
$\ll(\log x)^{2}q^{\frac{5}{16}}q^{\frac{5}{16}}(Mq)^{\frac{1}{6}}(Nq)^{\frac{1}{6}}L^{\frac{1}{2}}$
$\ll q^{\frac{3}{4}+\varepsilon}(MN)^{\frac{1}{16}}L^{\frac{1}{2}}$
@@
$T_{4}^{-}$ $\ll\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{2}\Big{)}^{\frac{7}{16}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{M}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{N}|^{6}\Big{)}^{\frac{1}{48}}\Big{(}\sum_{\begin{subarray}{c}\chi\in\Gamma_{4}^{-}\end{subarray}}|\mathcal{L}|^{12}\Big{)}^{\frac{1}{12}}$
$\ll(\log x)^{2}q^{\frac{7}{16}}q^{\frac{7}{16}}(Mq)^{\frac{1}{48}}(Nq)^{\frac{1}{48}}(qL^{2})^{\frac{1}{12}}$
$\ll q^{1+\varepsilon}(MN)^{\frac{1}{48}}L^{\frac{1}{6}}.$
@@
$\min\{S_{4}^{+},~{}T_{4}^{+}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,q^{-\frac{1}{4}}L^{\frac{1}{2}}\}\ll x^{\frac{1}{16}}q^{\frac{31}{32}+\varepsilon}$
@@
$\min\{S_{4}^{-},~{}T_{4}^{-}\}\ll q^{1+\varepsilon}(MN)^{\frac{1}{16}}\min\{1,L^{\frac{1}{6}}(MN)^{-\frac{1}{24}}\}\ll x^{\frac{1}{20}}q^{1+\varepsilon}.$
@@
Combining the above treatments in four parts, as well as the restrictions (5.18), (5.20) and (5.21), we choose
$M=N=x^{1-\varepsilon}q^{-\frac{6}{5}}$
@@
$D=\min\{x^{1-\varepsilon},~{}x^{2-\varepsilon}q^{-\frac{12}{5}}\}$
@@
Keep in mind that $q\geqslant x^{\frac{1}{2}}$ throughout this section, and we appeal to Proposition 4.1 with the level $D$ as large as possible.